### v1

# GTPBD: A Fine-Grained Global Terraced Parcel and Boundary Dataset

Abstract
Agricultural parcels serve as basic units for conducting agricultural practices and applications, which is vital for land ownership registration, food security assessment, soil erosion monitoring, etc. However, existing agriculture parcel extraction studies only focus on mid-resolution mapping or regular plain farmlands while lacking representation of complex terraced terrains due to the demands of precision agriculture. In this paper, we introduce a more fine-grained terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. GTPBD comprises 47,537 high-resolution images with three-level labels, including pixel-level boundary labels, mask labels, and parcel labels. It covers seven major geographic zones in China and transcontinental climatic regions around the world. Compared to the existing datasets, the GTPBD dataset brings considerable challenges due to the: (1) terrain diversity; (2) complex and irregular parcel objects; and (3) multiple domain styles. Our proposed GTPBD dataset is suitable for four different tasks, including semantic segmentation, edge detection, terraced parcel extraction and unsupervised domain adaptation (UDA) tasks. Accordingly, we benchmark the GTPBD dataset on eight semantic segmentation methods, four edge extraction methods, three parcel extraction methods and five UDA methods, along with a multi-dimensional evaluation framework integrating pixel-level and object-level metrics. GTPBD fills a critical gap in terraced remote sensing research, providing a basic infrastructure for fine-grained agricultural terrain analysis and cross-scenario knowledge transfer. The code and data are available at https://github.com/Z-ZW-WXQ/GTPBG/.
1 Introduction
In ancient China, terraced fields were given the reputation of "Dragon Spines" ("{CJK}UTF8gbsn龙脊" in Chinese), which is not only a poetic description of their winding and magnificent forms, but also a high praise in agricultural civilization. Nowadays, approximately 120 million acres of terraced fields worldwide still support more than 500 million mountainous populations unicef2024state , reducing soil erosion by more than 23.7 billion tons annually, which indicates significant ecological and economic values ferro2017terrace . Therefore, a better and fine-grained understanding of terraced area is needed for a sustainable agricultural ecosystem, which is vital for land ownership registration, food security assessment, etc spencer1961origin ; arnaez2015effects .
With the development of artificial intelligence techniques, terrace mapping have progressed significantly, from pixel-based and object-oriented methods to machine learning and deep learning-based methods from remote sensing images. Existing terrace mapping only focus on mid-resolution mapping cao202030 ; li202510 ; ding2021evaluation using Sentinel-1/2, Landsat, MODIS, etc. Meanwhile, some researchers have collected public large-scale agricultural parcels and boundaries datasets from high-resolution remote sensing images. To improve the agricultural parcel extraction, advanced deep learning methods have been used in parcel-level extraction and vectorization yu2022research ; pan2023e2evap ; ReauNet-Sober ; SEANet . Despite the impressive capabilities demonstrated by deep learning for agricultural parcel extraction, fine-grained terraced parcel and boundary benchmarks for agricultural remote sensing remain scarce due to they are extremely complex and unexpectedly irregular. Existing agricultural parcel dataset primarily focus on scenarios of regular plain farmlands (See Fig. 1 (a)). Datasets designed for agricultural parcel dataset from high-resolution remote sensing images exhibit notable limitations, primarily in terms of terrain diversity, annotation difficulty and multiple tasks, as shown in Fig. 1.
To begin with, current parcel dataset suffer from regular plain farmlands and limited terraced fields pan2023e2evap ; AI4Boundaries , limiting their ability to insufficient terrain diversity (See Fig. 1 (a)). To achieve a more precise agricultural mapping and yield estimation for global terraced field, it is imperative to develop a more fine-grained terraced dataset covering terrain diversity that contains stepped 3D structures, shadow-occluded scenes and irregular boundaries that combine three-dimensional topological structures (See Fig. 1 (b)). In addition, most of existing related datasets only provide binary classification mask labels FHAPD/HBGNet ; AI4Boundaries , failing to explicitly distinguish between two topological relationships of adjacent terraced ridges: shared edges (where adjacent parcels share the same boundary) and non-shared edges (where adjacent parcels have separate boundaries). It is highly demanded to collect a refined terraced parcel dataset with fine-grained boundaries. Furthermore, existing agricultural parcel dataset often rely on oversimplified task design such as binary classification or semantic segmentation Eurocropsml ; India , and have not discussed the transferability of agricultural parcel extraction. In order to improve and benchmark their generalizability, appropriate dataset that includes multiple domain styles is required.
Figure 1: Comparisons between previous dataset that focus on regular plain farmlands with limited terraced fields, and our GTPBD that offers more challenging and comprehensive evaluations.
In response to the aforementioned issues, this paper proposes a more challenging terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. In summary, the contributions of this work can be concluded as:
$\bullet$ We construct GTPBD which is the first global terraced dataset on parcel level, containing 47,537 images and collectively covering 885 $km^{2}$ of terraced farmland across varied terrains, with mountainous terraces accounting for over 80% of the dataset. The dataset covers seven major geographic zones in China and transcontinental climatic regions around the world.
$\bullet$ We propose GTPBD, which contains three-level labels for each pixel, including boundary labels, mask labels, and parcel labels. GTPBD is suitable for four different tasks, including semantic segmentation, edge detection, terraced parcel extraction and unsupervised domain adaptation (UDA) tasks with three different domains and six transfer tasks.
$\bullet$ We conduct a comprehensive evaluations of 8 semantic segmentation methods, 4 edge detection methods, 3 parcel extraction methods and 5 UDA methods, long with a multi-dimensional evaluation framework integrating pixel-level and object-level metrics. Experimental results reveal the limitations of current deep learning methods in terraced parcel extraction and its generalization capacity.
Table 1: Comparison between GTPBD and the main agricultural field datasets, including raster mask datasets and vector-only datasets.
Dataset Source Images Area (km2) Resolution Classes Global Task (m) coverage SS APE ED UDA Raster mask FHAPD FHAPD/HBGNet GaoFen-1/2 68,982 $<$500 0.5 2 FTW FTW Sentinel-2 70,462 166,293 10 2 AI4B (S-2) AI4Boundaries Sentinel-2 7,831 51,321 10 2 AI4B (Ortho) AI4Boundaries Aerial 7,598 1,992 1 2 PASTIS PASTIS Sentinel-2 2,433 3,986 10 19 CP-Set cp-set GF-2 949 249 1 2 GFSAD30 GFSAD30 Landsat-8 64,800 18,740,000 30 2 Vector-only GloCAB GloCAB Sentinel-2 190,832 N/A 10 2 EuroCropsML Eurocropsml Sentinel-2 706,683 >4,000 10 176 AI4SmallFarms AI4SmallFarms Sentinel-2 62 1,550 10 2 India Smallholder Airbus SPOT-6/7 10,000 30-50 1.5-4.8 2 Boundaries India & PlanetScope GTPBD (Ours) GF-2 & Google Earth 47,537 885 0.1–1 3
  • • SS: Semantic Segmentation; APE: Agricultural Parcel Extraction; ED: Edge Detection; UDA: Unsupervised Domain Adaptation.
2 Related works
2.1 Agricultural parcel and boundary datasets
In recent years, significant efforts have been made to establish comprehensive agricultural parcel datasets (see Table 1). Early datasets such as GFSAD30 GFSAD30 offer extensive global coverage but at a relatively coarse resolution of 30 meters, limiting their suitability for detailed parcel delineation tasks. PASTIS PASTIS and FTW FTW developed from Sentinel-2 imagery with resolutions around 10 meters, advanced the granularity of available agricultural data. FHAPD FHAPD/HBGNet and CP-Set cp-set comprise high-resolution GaoFen-2 (GF-2) imagery at 0.5 and 1 meters resolution, enhancing the spatial detail for agricultural parcel extraction tasks. But they predominantly include regularized, planar agricultural fields, limiting their applicability to terraced landscapes. AI4Boundaries AI4Boundaries provides annotations suitable for tasks such as semantic segmentation and edge detection, yet these efforts still primarily target flat agricultural areas, with limited consideration of terraced agricultural systems. Besides, vector-only datasets such as GloCAB GloCAB and EuroCropsML Eurocropsml though comprising a large number of images, still fall short of achieving true global coverage, and their utility is restricted to agricultural parcel extraction (APE) tasks because they only provide vector annotations. India Smallholder Boundaries India despite offering relatively high spatial resolution, focuses on flat farmlands in India and supports only APE tasks, making it less suitable for broader situation.
To overcome these limitations, we introduce a more challenging terraced parcel dataset named GTPBD, specifically designed to address the gaps in existing agricultural datasets. GTPBD is distinguished by its global scope, encompassing diverse regions spanning tropical, subtropical, and temperate climatic zones. Furthermore, GTPBD supports multiple tasks, including semantic segmentation, agricultural parcel extraction, unsupervised domain adaptation, and edge detection, offering a unified benchmark for advancing the analysis of complex terraced agricultural landscapes.
2.2 Remote sensing tasks
Semantic Segmentation (SS) assigns a class label to every pixel in imagery. Early encoder–decoder CNNs such as U-Net unet , DeepLabV3 deeplabv3 , and PSPNet pspnet , recover spatial detail via skip connections, atrous convolutions, and pyramid pooling. More recently, transformer-based frameworks including SegFormer segformer and Mask2Former mask2former , leverage self-attention to capture long-range dependencies and improve robustness to local appearance variations mask2former ; segformer ; transunet ; setr . Despite their advances, these models have been benchmarked almost on coarse-grained land-cover datasets FHAPD/HBGNet ; landcovernet ; GID , which focus on urban scenes or flat fields. To fill this gap, we evaluate all of the above architectures on GTPBD, which includes fine-grained annotations, to deliver the comprehensive semantic segmentation results for complex terraced agricultural landscapes.
Agricultural Parcel Extraction (APE) delineates farmland units for precision agriculture APEusage1 ; APEusage2 . Most CNN-based methods treated APE as single-task segmentation, using encoder–decoder backbones to predict parcel masks but often yielding unclosed or jagged boundaries in heterogeneous fields APEdisadvantage . To address this, edge-aware multi-task architectures have emerged: SEANet SEANet jointly predicts masks, distances, and edges to enforce closed parcels, REAUNet ReauNet-Sober integrates Sobel filters and attention for multi-scale edge enhancement, HBGNet FHAPD/HBGNet extracts low-level boundary features and high-level parcel semantics in parallel with a dual-branch framework. Despite these models perform well on planar benchmarks like FHAPD FHAPD/HBGNet and AI4Boundaries AI4Boundaries , their robustness to terraced ridge geometries is untested. We therefore benchmark them on GTPBD to provide systematic evaluation of APE methods in complex terraced agricultural landscapes.
Edge Detection (ED) precisely localizes pixel-level boundaries and facilitates high-fidelity terrace morphology analysis and erosion monitoring edgedetection1 ; edgedetection2 . We benchmark four state-of-the-art models on GTPBD: REAUNet-Sober ReauNet-Sober , which embeds Sobel-style filters into a U-Net backbone; MuGE Muge , employing multi-scale gating units to enhance edge features; PiDiNet pidinet , a lightweight pixel-difference network optimized for real-time inference; and UEAD UEAD , leveraging unsupervised edge-aware representation learning. Although these models have demonstrated strong performance on urban or synthetic benchmarks, their efficacy on complex terraced agricultural landscapes has not been evaluated. Our results on GTPBD’s fine-grained boundary annotations therefore provide accurate assessment of edge detection performance in terraced farmland environments.
Unsupervised Domain Adaptation (UDA) aims to transfer knowledge from a labeled source domain to an unlabeled target domain UDA1 . UDA methods for remote sensing can be broadly grouped into adversarial training and self-training approaches. Adversarial training such as Cycada cycada and CLAN CLAN leverage discriminators to align feature distributions, they incur complex optimization overhead and may not directly address style discrepancies. Representative self-training frameworks include DAFormer DAFormer , which integrates rare-class sampling and pseudo-label denoising. In addition, input-level adaptation methods like FDA FDA perform only low-frequency transfer in the Fourier domain, requiring no adversarial network or pseudo label. Although these UDA methods have been widely used in remote sensing community with some improvements, existing public remote sensing datasets have been limited by urban scenes li2024hyunida ; chen2025ban or land cover mapping wang2loveda ; zheng2024open . To this end, the GTPBD dataset is proposed for a more challenging benchmark, promoting future research of cross-regional or temporal agricultural and terraced parcel extraction algorithms and its applications.
Figure 2: The characteristics of our proposed GTPBD with diverse terrain, fine-grained annotation and multi-task support. GTPBD covers seven major geographic zones in China and transcontinental climatic regions around the world.
3 Dataset
3.1 Image collection and manual annotation
We collect our images in our dataset based on the Global Terrace Map (GTM) with 10-m resolution li202510 . To ensure data quality and spatial coverage, we adopted the following screening strategies:
Spatial Distribution We give selection priority to regions with dense terrace distribution and high topographic complexity according to GTM li202510 . We also select typical terraces in diverse landforms, such as the alpine terraces in southwestern China and the Sa Pa terraces in Vietnam, to evaluate the model’s generalization ability across varied landforms. Fig 2 displays the major study areas covering fourteen countries around the world, including China, Vietnam, Tunisia, Ethiopia, Peru, Mexico and more. In addition, our dataset covers seven major geographic regions of China, including Central China, South China, North China, East China, Northwest China, Northeast China, and Southwest China. More images could be found in Appendix B.1.
Image Resolution We download a series of high-resolution terraced field images from Google Earth, with 0.1$\sim$1m, ensuring the delicacy and clarity of the images. More importantly, these high-resolution terraced field images not only cover diverse seasonal changes but also capture the scenery under different lighting conditions, thus comprehensively showcasing the appearance of terraced fields in different environments. It is worth mentioning that all of these images were carefully selected from cloud-free terraced fields between 2021 and 2025, providing extremely valuable and high-quality data resources for related research.
Parcel Annotation We construct the topological vectorization and annotations in terraced fields through QGIS. We recruit over 50 participants, comprising undergraduate and graduate students, to manually annotate the fine-grained terraced parcel and boundary with strict quality inspection. The specific implementation strategy is as follows: for areas with dense hierarchical structures and narrow boundaries (less than 0.5m), the common edge annotation method is adopted - prioritizing vectorization of large continuous parcels, and generating sub-parcel topologies through internal line feature cutting; For areas where the field ridge width is greater than or equal to 0.5 m, bilateral segmentation annotation is applied — independent vector boundaries are generated along both sides of the field ridge, and the ridge itself is removed to form a non-standard edge topology structure.
Three-Level Labels As for mask labels, we use GDAL’s rasterize function for fully connected rasterization (all touched strategy), establish a binary semantic segmentation benchmark, and assign 1 to terraced areas and 0 to backgrounds. As for boundary labels, we use a 3$\times$3 rectangular grid mask to perform a single morphological etching and construct a 3-pixel-wide edge detection dedicated label. As for parcel labels, we generate pure parcel regions through mask boundary using the XOR operation to achieve terraced parcel extraction tasks. Appendix B.2 displays the differences between previous dataset and our GTPBD.
Dataset Construction We crop the original remote sensing image with the size of 512$\times$512 pixels according to existing works AI4Boundaries . We divide the GTPBD dataset into training set, validation set, and testing set by the raito of 60%, 20% and 20% (see Fig. 3 (b)), which are used for model training, tuning and hyperparameter selection, and performance evaluation, respectively.
3.2 Statistics
Figure 3: Statistics for the number and area of GTPBD dataset. (a) Distribution of images across different regions. (b) Distribution of images across different dataset splits. (c) Distribution of area across different regions. (d) Distribution of the parcel sizes (logarithmic scale in vertical axis).
[…]
3.3 Differences among three domains
Figure 4: Statistics for three different domains. (a) Distribution of area across different domains. (b) The number of images across different domains. (c) Spectral statistics of mean and standard deviation ($\sigma$) for different domains. (d) Distribution of parcel sizes across different domains (logarithmic scale).
Figure 5: Some cases for three different domains: Global, North and South.
We seperate our GTPBD into three domains: South China (South), North China (North) and other regions except China (Global). The detailed statics of these three domains are shown in Fig. 4 (a) and (b). According to Fig. 5, the characteristics from different regions are really different. For the spectral statistics, the mean values are similar (Fig. 4 (c)) because of the large-scale homogeneous geographical areas and diverse land cover types. We could observe that the South domain have lower standard deviations. As is shown in Fig. 4 (d), although all domains exist “long tail” phenomenon (logarithmic scale in vertical axis), the South domain has the most of the parcels have relatively small scales. When faced with large-scale terraced parcel and boundary extraction tasks, the differences between different scenes or regions bring new challenges to the model generalization and transferability.
4 Experiments
4.1 Experimental setup
The data splits followed Sec. 3.1. During the training, we used the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay of $10^{-4}$. For the data augmentation, 512$\times$512 patches were randomly cropped with random mirroring and rotation. We implemented all our experiments on NVIDIA RTX 4090 GPU. We benchmark the GTPBD dataset on eight semantic segmentation (U-Net unet , DeepLabV3 deeplabv3 , PSPNet pspnet , NonLocal nonlocal , OCRNet OCRNet , K-Net knet , SegFormer segformer and Mask2Former mask2former ) methods, four edge extraction methods (MuGE Muge , PiDiNet pidinet , UEAD UEAD and REAUNet-Sober ReauNet-Sober ), three parcel extraction methods (REAUNet ReauNet-Sober , SEANet SEANet , HBGNet FHAPD/HBGNet ) and five UDA methods (Source only, FDA FDA , PiPa PiPa , HRDA HRDA and DAFormer DAFormer ). More implementation details are provided in the Appendix C.
4.2 Evaluation metrics
We brief list our comprehensive evaluation metrics for different tasks including pixel-level, object-level and edge detection of GTPBD. More details could be found in Appendix D.
(1) Pixel-level evaluation metrics: To evaluate segmentation accuracy on the GTPBD dataset, we adopt five standard pixel-level metrics commonly used in semantic segmentation and unsupervised domain adaptation (UDA) tasks, including Precision (Prec.), Recall (Rec.), Intersection over Union (IoU), Overall Accuracy (OA) and F1-score; (2) Object-level geometric metrics: To evaluate the geometric accuracy of parcel delineation on the GTPBD dataset, we adopt three object-level metrics: Global Over-Classification Error (GOC), Global Under-Classification Error (GUC), and Global Total Classification Error (GTC). These indicators provide a comprehensive assessment of shape fidelity and segmentation precision at the object level; (3) Edge detection evaluation metrics: For edge detection tasks on the GTPBD dataset, we evaluate model performance using three standard metrics: Optimal Dataset Scale F1-score (ODS), Optimal Image Scale F1-score (OIS), and Average Precision (AP). These metrics jointly assess the adaptability and stability of boundary predictions.
Table 2: Comparisons of different semantic segmentation methods on the GTPBD dataset (%). More visualization results and the performance of regions could be found in Appendix E.1.2 and E.1.1.
[…]
U-Net unet 78.21 72.79 60.50 74.18 75.40
Deeplabv3 deeplabv3 81.57 81.64 68.93 79.99 81.61
PSPNet pspnet 79.03 84.77 69.20 79.48 81.80
Nonlocal nonlocal 82.67 80.22 68.67 80.10 81.43
OCRNet OCRNet 77.53 86.22 68.98 78.92 81.65
K-Net knet 78.69 81.18 66.55 77.81 79.92
SegFormer segformer 80.48 83.03 69.11 79.82 81.73
Mask2Former mask2former 73.02 94.25 69.91 77.94 82.29
Table 3: Comprehensive performance comparison of edge detection models on the GTPBD dataset.
[…]
MuGE Muge 62.56 61.93 65.12
PiDiNet pidinet 53.70 53.12 52.92
UEAD UEAD 25.88 26.01 17.94
REAUNet-Sober ReauNet-Sober 65.06 63.73 70.09
4.3 Semantic Segmentation Results
Semantic segmentation plays a vital role in agricultural parcel extraction. As summarized in Table 3, these models exhibit notable discrepancies in segmentation performance, reflecting varying capabilities in modeling fine-grained agricultural boundaries. Specifically, the NonLocal model achieved the highest Prec. (82.67%) and OA (80.10%), indicating strong discrimination with minimal false positives. In contrast, Mask2Former demonstrated superior generalization, attaining the best scores in Rec. (94.25%), IoU (69.91%), and F1-score (82.29%). These findings highlight a trade-off between precision and recall among different architectures and emphasize the robustness of transformer-based models under complex terraced scenarios. Fig. 6 visually compare the segmentation results on a sample patch, with red highlighting false positives and blue highlighting false negatives. More visualization results and the performance of regions could be found in Appendix E.1.2 and E.1.1.
Figure 6: Qualitative comparisons of semantic segmentation and error maps among different methods.
4.4 Edge Detection Results
Conventional edge detection datasets typically include multi-scale, manually annotated boundaries. To simulate this for the GTPBD dataset, we apply morphological operations (erosion and dilation) to generate edge labels with widths of 1 to 5 pixels, mimicking human-labeled multi-resolution boundaries. Table 3 presents the results of four edge detection models. Among them, REAUNet-Sober achieves the best performance across all three metrics—ODS (0.6506), OIS (0.6373), and AP (0.7009)—significantly outperforming both general-purpose (e.g., PiDiNet), demonstrating its effectiveness in delineating parcel edges under noisy conditions. Visual examples in Fig. 7 (f) further illustrate model robustness, where false positives are marked in red and missed detections in blue.
Figure 7: Edge detection and terraced parcel extraction. (a) Probability Map. (b) Binarized Prediction. (c) Ground Truth. (d) Error Regions (Red: FP, Blue: FN). (e) Overlay Visualization (alpha=0.5). (f) The performance of integrating edge results (REAUNET-Sober) and semantic segmentation results. (g) Evaluation of object-level parcel extraction (Red: GOC, Blue: GUC, Green: GT boundary).
Table 4: Performance Comparison of Parcels Extraction(unit:%)
[…]
U-Net 83.27 75.69 65.22 79.87 78.95 66.11 37.33 38.80
PSPNet 83.70 80.56 69.64 82.23 82.10 64.75 38.75 40.15
SegFormer 82.45 82.19 69.95 82.24 82.32 63.39 39.85 40.87
Mask2Former 82.59 81.83 69.80 82.19 82.21 62.98 39.40 40.55
REAUNEet 78.44 76.79 63.41 78.98 77.61 23.77 53.05 25.18
SEANet 69.85 88.58 64.08 76.54 78.11 11.50 71.96 14.82
HGBNet 76.14 87.82 68.87 81.17 81.56 8.58 67.39 12.22
Table 5: Comparisons of UDA and Boundary-integration UDA Performance on GTPBD (%).
Conventional UDA Boundary-integration UDA (Integrated by REAUNet-Sober)
[…]
S $\rightarrow$ N Source Only 75.93 82.12 70.05 77.47 78.92 73.11 83.21 68.36 76.71 77.87 8.94 49.80 11.68
FDAFDA 82.60 78.20 67.13 79.27 80.34 79.55 77.95 64.94 78.26 78.74 25.04 46.09 24.44
PiPa PiPa 88.40 81.43 73.57 84.16 84.77 88.40 81.43 73.57 84.16 84.77 24.59 16.63 11.90
HRDA HRDA 88.00 82.44 74.11 84.48 85.13 84.00 83.44 72.00 83.24 83.72 18.89 41.22 19.87
DAFormer DAFormer 87.23 82.71 73.78 84.08 84.91 84.83 83.63 72.75 83.72 84.23 17.52 39.94 19.22
S $\rightarrow$ G Source Only 74.66 68.19 60.69 72.16 75.54 71.54 70.14 60.53 72.44 75.41 24.59 49.02 23.90
FDAFDA 76.27 87.16 68.57 74.78 81.35 81.18 75.68 63.38 74.78 78.33 22.11 48.85 22.81
PiPa PiPa 83.90 75.22 65.73 75.25 79.32 83.90 75.22 65.73 75.25 79.32 29.28 21.09 12.94
HRDA HRDA 61.35 95.53 59.64 66.31 74.71 78.22 77.30 63.61 73.36 77.76 20.93 50.11 21.94
DAFormer DAFormer 82.89 80.27 68.87 77.10 81.56 79.64 82.05 67.82 76.55 80.82 17.53 49.80 21.01
N $\rightarrow$ S Source Only 64.35 80.32 55.58 65.93 61.45 54.97 78.98 47.95 63.27 64.82 13.83 73.75 19.95
FDAFDA 67.40 71.73 53.25 67.20 69.50 59.08 68.04 46.25 66.11 63.25 29.48 60.68 33.65
PiPa PiPa 77.68 61.11 51.98 70.59 68.40 77.68 61.11 51.98 70.59 68.40 42.26 30.86 26.13
HRDA HRDA 76.90 60.10 50.91 69.81 67.47 66.52 65.74 49.39 71.14 66.13 32.03 58.84 36.01
DAFormer DAFormer 79.64 60.69 52.54 71.44 68.89 68.68 68.03 51.92 73.01 68.35 25.72 62.58 33.00
N $\rightarrow$ G Source Only 72.94 72.51 63.10 73.27 77.38 70.05 74.88 63.10 73.62 77.38 21.50 51.35 22.35
FDAFDA 84.17 74.95 65.69 75.30 79.29 81.18 75.68 63.38 74.78 78.33 22.11 48.85 22.81
PiPa PiPa 83.52 76.03 66.11 75.40 79.60 83.52 76.03 66.11 75.40 79.60 27.74 22.01 13.36
HRDA HRDA 87.45 66.81 60.96 73.01 75.75 84.59 68.04 60.54 73.28 75.42 27.13 42.88 24.07
DAFormer DAFormer 83.56 77.49 67.24 76.17 80.41 79.98 78.75 65.79 75.32 79.36 18.40 49.54 21.72
G $\rightarrow$ N Source Only 76.82 80.44 69.52 77.30 78.59 74.21 81.99 68.46 76.96 77.93 9.71 48.20 12.46
FDAFDA 71.77 89.22 66.04 75.16 79.55 72.10 82.14 62.33 74.35 76.79 21.83 54.51 24.76
PiPa PiPa 74.40 92.08 69.93 78.56 82.30 74.40 92.08 69.93 78.56 82.30 10.00 34.23 9.47
HRDA HRDA 75.42 91.37 70.41 79.20 82.63 72.67 90.95 67.77 77.66 80.79 9.33 52.11 13.88
DAFormer DAFormer 85.40 79.72 70.16 81.63 82.46 81.99 79.93 67.99 80.56 80.95 19.52 44.51 20.39
G $\rightarrow$ S Source Only 54.12 71.52 45.98 55.97 61.78 54.65 80.53 48.27 63.02 65.11 13.66 74.62 20.22
FDAFDA 59.39 88.36 55.09 62.47 71.04 51.21 81.85 45.99 58.81 63.01 21.59 66.56 28.27
PiPa PiPa 65.25 88.79 60.29 69.53 75.22 65.25 88.79 60.29 69.53 75.22 16.14 32.04 12.58
HRDA HRDA 61.35 95.53 59.64 66.31 74.71 66.30 95.09 60.33 70.05 75.33 16.20 66.58 28.68
DAFormer DAFormer 70.04 91.15 63.72 70.60 77.11 67.77 91.66 61.91 70.65 75.18 14.32 64.13 26.83
Avg Source Only 69.14 75.52 60.49 70.01 72.61 66.43 76.82 59.11 69.52 73.76 15.87 56.93 18.10
FDAFDA 73.60 81.60 61.96 72.53 77.18 70.55 76.89 58.38 72.38 75.38 23.36 54.09 25.46
PiPa PiPa 78.86 79.77 64.94 75.91 78.27 80.06 77.78 65.61 77.87 78.47 24.00 26.81 14.73
HRDA HRDA 78.18 81.05 62.28 74.52 78.40 75.72 77.43 60.86 76.79 76.53 24.12 51.79 24.41
DAFormer DAFormer 81.13 78.67 66.39 76.84 79.56 77.15 77.84 64.82 76.63 78.16 19.58 53.90 23.36
4.5 Boundary-integration terraced parcel extraction results
Building on the boundary constraint theory by Yuan et al. Yuhui2020 , which emphasizes the role of edge orientation and intensity in enhancing object-level delineation, we explore boundary-integration strategies for agricultural parcel segmentation. We compare four general-purpose segmentation models (U-Net, PSPNet, SegFormer, Mask2Former) with three agriculture-specific ones (HBGNet FHAPD/HBGNet , REAUNet ReauNet-Sober , SEANet SEANet ). To incorporate edge cues, we introduce a boundary detection pipeline. REAUNet-Sober ReauNet-Sober achieves the best performance among trained edge detectors, and its binarized output is integrated into the segmentation process. As shown in Table 4, REAUNet ReauNet-Sober delivers the best object-level performance with the lowest GUC with 53.05%, despite lower pixel-level scores. In contrast, SegFormer segformer attains the highest pixel-level accuracy (OA: 82.24%) but suffers from higher object-level errors, indicating a trade-off between pixel precision and structural integrity.
4.6 Unsupervised domain adaptation results
We evaluate domain robustness of segmentation models through UDA experiments across three distinct domains in GTPBD (see Sec. 3.3). The Source only and other four domain adaptation methods are benchmarked: FDA, DAFormer, HRDA, and PiPa. As shown in Table 5, under standard (non-boundary-integration) settings, DAFormer DAFormer achieves the highest average performance across Prec., IoU, OA, and F1, demonstrating strong generalization under domain shift. After adding boundary results (integrated by the performance of REAUNet-Sober ReauNet-Sober ), PiPa PiPa consistently outperforms others, achieving top scores across both pixel-level (Prec., IoU, OA, F1) and object-level metrics, including the lowest GUC and GTC. This highlights its advantage in preserving parcel structure with fine-grained segmentation. Visualization comparisons are provided in Appendix E.4.
5 Conclusion
In this paper, we propose a more challenging terraced parcel dataset named GTPBD (Global Terraced Parcel and Boundary Dataset), which is the first fine-grained dataset covering major worldwide terraced regions with more than 200,000 complex terraced parcels with manually annotation. GTPBD comprises 47,537 high-resolution images with three-level labels, including pixel-level boundary labels, mask labels, and parcel labels. It covers seven major geographic zones in China and transcontinental climatic regions around the world. Our proposed GTPBD dataset is suitable for four different tasks, including semantic segmentation, edge detection, terraced parcel extraction and Unsupervised Domain Adaptation (UDA) tasks. Accordingly, we benchmark the GTPBD dataset on eight semantic segmentation methods, four edge extraction methods, three parcel extraction methods and five UDA methods, along with a multi-dimensional evaluation framework integrating pixel-level and object-level metrics. Results highlight the limitations of current models in extracting fine-grained terraced parcels and boundaries, especially under domain shifts. Therefore, our proposed GTPBD fills a critical gap in remote sensing benchmarks and provides a critical infrastructure for complex agricultural terrain analysis and cross-scenario knowledge transfer.
Appendix A Broader impact
The proposed GTPBD dataset provides the first large-scale, fine-grained benchmark for terraced parcel analysis across diverse global terrains. Its release is intended to facilitate progress in computer vision methods for agricultural mapping, with potential societal benefits including improved food security analysis, precision agriculture, and climate-adaptive land management. By focusing on underrepresented mountainous regions, GTPBD promotes algorithmic development that extends beyond conventional flat farmland, contributing to more inclusive and geographically equitable AI research.This dataset may assist governmental and environmental agencies in monitoring land use, identifying erosion risks, and planning sustainable agricultural practices. It may also support digital infrastructure for land ownership registration in developing regions, where documentation remains limited or inaccessible.However, the deployment of models trained on GTPBD in real-world scenarios must be approached with care. Potential negative impacts include misuse of parcel delineation algorithms for land commodification, surveillance, or disempowerment of local communities, especially in regions with contested or informal land tenure. Furthermore, any bias introduced during data annotation or model training could disproportionately affect underrepresented regions or farming systems.
To mitigate these risks, we strongly encourage practitioners to collaborate with local stakeholders and domain experts, ensure transparency in model usage, and align applications with ethical land governance principles. The dataset should serve as a scientific tool for advancing equitable and sustainable land analysis, not for enabling exploitative practices.
Appendix B More image cases
B.1 More image cases in GTPBD
Figure 8: More images cases of different regions in GTPBD.
GTPBD covers seven major geographic zones in China and transcontinental climatic regions around the world. There are different terraces in different regions, and Fig. 8 supplements the images of terraces of the 12 regions mentioned in Fig. 2 .
B.2 Some cases in previous datasets
Fig. 9 displays typical annotation–imagery pairs from six widely used parcel datasets: (a) FHAPD [48], (b) FTW [17], (c) AI4Boundaries [9], (d) PASTIS [12], (e) CP-Set [46], and (f) GFSAD30 [31]. Although each benchmark has advanced parcel delineation research, several common limitations are evident. FHAPD and CP-Set rely on binary masks with coarse edges, leaving thin ridges either over-smoothed or missing. FTW and PASTIS are built on Sentinel-2 imagery (10 m GSD), so images appear blurry and smallholder fields are merged. AI4Boundaries depicts large, orthogonal parcels from mechanised farms, providing little shape diversity. GFSAD30, at 30 m resolution, captures only blocky field outlines and omits fine-scale geometry altogether. These issues—low spatial resolution, overly regular parcel shapes, and imprecise or incomplete boundaries limit the evaluation of models on fragmented, irregular terrains.
Figure 9: Representative Samples from Existing Agricultural Parcel Datasets
B.3 More statistics of GTPBD
Figure 10: Distribution of parcel sizes across different regions. (logarithmic scale)
In Sec. 3.2, detailed statistical analyses of the data have been carried out. To clearly illustrate the distribution of parcel sizes across the eight regions, Fig. 10 provides individual bar charts for each area. These charts highlight that small-sized terraced blocks dominate in every region.
Appendix C Model details
C.1 Semantic segmentation
MMSegmentation is an open-source semantic segmentation toolbox built on the OpenMMLab ecosystem and PyTorch  [^1] https://github.com/open-mmlab/mmsegmentation. It implements a wide range of state-of-the-art architectures, such as U-Net, DeepLabV3, PSPNet, and Segformer, thereby facilitating reproducible comparisons across models. Our semantic segmentation experiments are all completed under the MMsegmention framework.
C.1.1 U-Net
U-Net, proposed by Ronneberger et al. [28], adopts a symmetric encoder–decoder design that captures contextual information via a contracting path and enables precise localization through an expansive path. In our experiments using MMSegmentation’s default U-Net implementation, the key hyperparameters are configured as Table 6.
Table 6: The hyperparameters of U-Net
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone base channels Number of filters in the first convolution layer 64 Number of stages Levels of encoder–decoder (depth of U-Net) 5 Convolutions per stage Number of conv layers in each block 2 Optimizer Optimization algorithm and initial learning rate Adam, lr=$1e^{-4}$ Sliding-window crop size Sliding-window inference patch size 256 × 256 Sliding-window stride Stride between adjacent sliding-window patches 170 Training iterations Total number of training iterations 20,000
C.1.2 DeepLabV3
DeepLabV3 [6] enhances semantic segmentation by employing atrous (dilated) convolutions to enlarge the receptive field without losing resolution, and by integrating an Atrous Spatial Pyramid Pooling (ASPP) module to capture multi-scale context. In our experiment, All hyper-parameters remain the same as in the mmsegmentation framework and are summarized in Table 7.
Table 7: The hyperparameters of DeeplabV3
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Type and depth of backbone network ResNetV1c-50 Backbone dilations Dilation rates for each backbone stage (1,1,2,4) Backbone strides Stride sizes for each backbone stage (1,2,1,1) ASPP dilations Dilation rates in Atrous Spatial Pyramid Pool (1,12,24,36) ASPP channels Number of intermediate channels in ASPP head 512 Auxiliary head channels Number of channels in auxiliary FCN head 256 Number of classes Number of segmentation target classes 2 Training iterations Total number of training iterations 20,000
C.1.3 PSPNet
PSPNet [49] introduces a Pyramid Pooling Module that aggregates context information at multiple spatial scales by pooling features into different bin sizes, and then upsampling and concatenating them with the original feature map. This design enables the network to capture both global context and fine details simultaneously. In our experiments, we rely on the out-of-the-box configuration provided by MMSegmentation without further modification, and summarize the key hyperparameters in Table 8.
Table 8: The hyperparameters of PSPNet
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Type and depth of backbone network ResNetV1c-50 Backbone dilations Dilation rates for each backbone stage (1,1,2,4) Backbone stages Number of residual stages used 4 PSP pooling scales Sizes of the pyramid pooling bins (1,2,3,6) PSP intermediate channels Number of feature channels after pooling 512 Auxiliary head channels Number of channels in auxiliary FCN head 256 Loss weights Main and auxiliary segmentation loss weights 1.0 (main), 0.4 (auxiliary) Training iterations Total number of training iterations 20,000
C.1.4 NonLocal
Non-local Neural Networks [38], a generic non-local operation that computes pairwise interactions between any two positions on the feature map, enabling the model to capture long-range dependencies and context beyond local receptive fields. We adopt the standard MMSegmentation implementation without modifying its default settings and report the core hyperparameters in Table 9.
Table 9: The hyperparameters of NonLocal
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Type and depth of backbone network ResNetV1c-50 Backbone dilations Dilation rates for each backbone stage (1,1,2,4) NLhead intermediate channels Number of feature channels in the NLHead 512 NL reduction ratio Channel reduction factor inside the non-local block 2 NL operation mode Similarity function used in non-local computation embedded_gaussian Dropout ratio Dropout probability applied in the NLHead 0.1 Training iterations Total number of training iterations 20,000
C.1.5 OCRNet
OCRNet [43] enhances pixel-wise segmentation by introducing an object-contextual representation module that aggregates contextual information from object regions to refine per-pixel predictions. Specifically, it first generates a coarse object region map via soft object-region pooling, then computes a pixel-wise contextual representation by weighting features according to their belongingness to these regions. This two-stage design allows the network to capture both local details and global object semantics effectively. Table 10 summarize the main hyperparameters in our experiment.
Table 10: The hyperparameters of OCRNet
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Type and depth of backbone network HRNetV2-W18 Number of cascade stages Number of encoder–decoder stages 2 FCNhead intermediate channels Total channels after concatenating multi-scale features 270 OCRHead channels Number of feature channels in OCR contextual module 512 OCR channels Reduced channel dimension inside OCR module 256 Training iterations Total number of training iterations 20,000
C.1.6 K-Net
K-Net [47] formulates semantic segmentation as a dynamic kernel-based grouping problem, enabling adaptive context aggregation and object-level reasoning. Each kernel generates an attention map that segments the feature map into distinct semantic regions, and these kernels are updated via a lightweight transformer-style interaction. In our experiments, we employ MMSegmentation’s K-Net configuration without further modifications and summarize the core hyperparameters in Table 11.
Table 11: The hyperparameters of K-Net
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Type and depth of backbone network ResNetV1c-50 Decode head stages Number of iterative decoding stages 3 Kernel update FFN channels Hidden dimension in feed-forward networks of update heads 2048 Kernel update attention heads Number of attention heads in each KernelUpdateHead 8 Convolution kernel size Kernel size for pointwise conv in update heads 1 Training iterations Total number of training iterations 20,000
C.1.7 Segformer
SegFormer [39] leverages a hierarchy of lightweight MLP-based Mix-FFN blocks and multi-level feature fusion to achieve efficient and effective semantic segmentation without positional encodings. Its simple encoder–decoder design employs a series of Transformer-like layers that progressively downsample the input, and fuse multi-scale representations via a lightweight MLP decoder. Key parameters are shown in Table 12 in our experiment.
Table 12: The hyperparameters of Segformer
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone Type of hierarchical Transformer encoder MixVisionTransformer Embed dims Dimension of the token embeddings at stage 1 32 Number of stages Levels of hierarchical feature extraction 4 Layers per stage Number of attention heads in each KernelUpdateHead 8 Layers per stage Transformer blocks in each stage [2,2,2,2] Attention heads Number of self-attention heads per stage [1,2,5,8] Patch sizes Size of convolutional patches at each stage [7, 3, 3, 3] SR ratios Spatial reduction ratios before attention [8,4,2,1] Training iterations Total number of training iterations 20,000
C.1.8 Mask2Former
Mask2Former [8] presents a universal architecture for segmentation tasks by modeling pixel-wise, instance-wise, and semantic-level grouping using a mask-based Transformer decoder. It decouples mask prediction from task-specific heads, employing dynamic queries and multi-scale features to produce segmentation masks across different granularities. We employ Mask2Former configuration directly without any additional tuning and the key hyperparameters can be seen in Table 13.
Table 13: The hyperparameters of mask2former
Hyperparameter Description Typical Value Input image size Dimensions of input images 512×512 Backbone architecture Feature extractor type and depth ResNet-50 Number of queries Number of learnable mask queries 100 Transformer encoder layers Number of layers in the Deformable DETR transformer encoder 6 Transformer decoder layers Number of layers in the Mask2Former transformer decoder 9 Classification loss weight Weight for the classification loss in mask prediction 2.0 Mask & Dice loss weights Weights for mask and Dice losses mask=5.0, dice=5.0 Training iterations Total number of training iterations 20,000
C.2 Edge Detection Methods
C.2.1 UEAD
UEAD [53] addresses the annotation ambiguity and subjectivity inherent in edge detection by modeling label uncertainty. Instead of relying on deterministic pixel-wise labels, it learns a Gaussian distribution over annotations to capture labeling variance. The estimated variance serves as a measure of pixel-level uncertainty, and an adaptive weighting loss emphasizes learning from uncertain (i.e., hard) samples. UEAD can be integrated with various encoder–decoder backbones and consistently improves performance across multiple edge detection benchmarks.
C.2.2 MuGE
MuGE[52] introduces a novel edge detection framework that captures edge ambiguity by generating edge maps at multiple controllable granularities. It first predicts edge granularity from annotations, then injects this granularity into multi-scale feature maps to produce edge maps ranging from coarse contours to fine textures. By decomposing feature maps into frequency components, MuGE enables fine control over edge detail, resulting in both interpretability and high accuracy.
C.2.3 PiDiNet
PiDiNet[30] offers a lightweight and efficient solution for edge detection by combining traditional gradient operators with modern convolutional design. It introduces pixel difference convolutions that emulate classical detectors like Canny and Sobel, enabling real-time inference with high accuracy. With less than 1M parameters, PiDiNet achieves human-level performance on BSDS500 and maintains strong efficiency–accuracy trade-offs across diverse benchmarks.
C.3 Agricultural Parcel Extraction Methods
C.3.1 REAUNet
REAUNet[21] is a tailored edge-aware network designed for accurate agricultural parcel delineation from both medium- and high-resolution remote sensing imagery (Sentinel-2 and GF-2). It addresses the common issue of unclosed and fragmented parcel edges by integrating four key components: an edge detection block, a dual attention block, a deep supervision mechanism, and a refine module. These modules collectively enhance the model’s sensitivity to boundary information and improve multi-scale consistency. Experiments show that all four components are critical, with full REAUNet outperforming partial variants. Compared with SEANet and MPSPNet, REAUNet achieves improvements in both thematic (F1, IoU) and geometric (GTC) accuracy, and demonstrates promising transferability to unseen regions with limited fine-tuning data.
C.3.2 SEANet
SEANet[18] is a multi-task neural network that simultaneously predicts semantic masks, edges, and distance maps for precise and closed parcel delineation. By explicitly modeling boundary extraction as an edge detection task, SEANet enhances geometric accuracy and is particularly effective for small, irregular parcels. It incorporates a multi-level edge feature extraction mechanism and a task uncertainty-aware loss to improve generalization. Extensive experiments on both high-resolution (GF-2) and medium-resolution (Sentinel-2) images across China and Europe show that SEANet produces more accurate parcel layouts and demonstrates robust cross-region transferability.
C.3.3 HBGNet
HBGNet[48] is a hierarchical dual-branch framework designed to fully exploit boundary semantics for robust agricultural parcel extraction. It consists of a core AP extraction branch and an auxiliary boundary branch enhanced by Laplacian-based convolution. To improve adaptability across varying parcel sizes and morphologies, HBGNet integrates global–local context aggregation and boundary-guided fusion modules. It also introduces FHAPD, the first large-scale VHR agricultural parcel dataset in China, to support comprehensive evaluation. HBGNet outperforms eight existing methods across multiple datasets (FHAPD, Al4Boundaries, Sentinel-2) in both attribute and geometric metrics, achieving up to 7.5.
Figure 11: Regional Segmentation Performance Across Models
Figure 12: Error Map Comparison of different Semantic Segmentation Models
C.4 Unsupervised Domain Adaptation Methods
C.4.1 FDA
FDA[41] is a simple yet effective UDA method that reduces domain gaps by aligning the low-frequency components of source and target images in the Fourier domain. The core idea is that low-frequency information in images (e.g., color and style) mainly accounts for the domain shift, while high-frequency components capture structural details relevant to semantic content. By replacing the low-frequency spectrum of source images with that of target images, FDA transfers global appearance cues without altering the semantic layout. This process improves feature-level alignment and enhances segmentation performance in the target domain without requiring architectural changes or additional supervision.
C.4.2 DAFormer
DAFormer[15] introduces a Transformer-based architecture for UDA in semantic segmentation, leveraging the superior representation capability of vision Transformers. The model integrates a Transformer encoder with a multi-level, context-aware decoder and employs three essential training strategies: rare class sampling to improve pseudo-labels, ImageNet feature distance regularization to stabilize transfer, and learning rate warmup to reduce overfitting. These components collectively enable DAFormer to effectively adapt to the target domain and learn rare or hard classes. DAFormer achieves significant improvements over previous UDA methods, setting new benchmarks on synthetic-to-real segmentation tasks.
C.4.3 HRDA
HRDA[16] proposes a multi-resolution UDA framework to balance fine-detail preservation and long-range context modeling in semantic segmentation. It combines small high-resolution image crops—effective for boundary and small object delineation—with large low-resolution crops that retain broader contextual information. A learned scale attention module adaptively fuses these dual-resolution representations. HRDA significantly improves segmentation quality on domain shift benchmarks while maintaining computational efficiency, and is particularly effective for detailed structures under limited GPU memory.
C.4.4 PiPa
PiPa[23] introduces a unified pixel- and patch-level self-supervised learning framework for UDA in semantic segmentation. Unlike traditional UDA methods that only minimize inter-domain discrepancies, PiPa focuses on enhancing intra-domain consistency by modeling pixel-level intra-class compactness and patch-level context-invariant semantics. This dual-level supervision enables the model to learn more robust and domain-invariant representations. PiPa achieves competitive results on standard UDA benchmarks and can be flexibly combined with other UDA methods to further improve performance without increasing model complexity.
Figure 13: Qualitative edge detection comparisons between PiDiNet [30] and UEAD [53].
Appendix D Evaluation metrics
D.1 Pixel-level evaluation metrics
To evaluate segmentation accuracy on the GTPBD dataset, we adopt five standard pixel-level metrics commonly used in semantic segmentation and unsupervised domain adaptation (UDA) tasks, including Precision (Prec.), Recall (Rec.), Intersection over Union (IoU), Overall Accuracy (OA) and F1-score.
Precision (Prec.) measures the proportion of correctly predicted agricultural pixels among all pixels predicted as agricultural. A higher precision indicates fewer false positives in land parcel classification.
$Prec.=\frac{TP}{TP+FP}\quad$ (1)
where TP and FP indicate true positive and false positive, respectively. TP indicates the number of pixels correctly identified as agricultural parcels, while FP indicate the number of pixels mis-identified as agricultural parcels (i.e., mistakes).
Recall (Rec.) captures the proportion of true agricultural pixels that are correctly identified. High recall reflects a model’s ability to minimize missed detections.
$Rec.=\frac{TP}{TP+FN}\quad$ (2)
where FN indicates false negative and the number of pixels mis-identified as non-agricultural parcels (i.e., omissions).
Intersection over Union (IoU) quantifies the spatial agreement between the predicted and ground-truth regions. IoU is widely adopted in segmentation benchmarks due to its robustness to class imbalance.
$IoU=\frac{TP}{TP+FP+FN}\quad$ (3)
Figure 14: Qualitative edge detection comparisons between REAUNet-Sober [21] and MuGE [52].
Figure 15: Regional Benchmark of Agricultural Parcel Extraction Models across Eight Metrics
Overall Accuracy (OA) calculates the proportion of correctly classified pixels across the entire image. OA provides a global view of model performance and is especially informative for datasets with class imbalance.
$OA=\frac{TP+TN}{TP+FP+FN+TN}\quad$ (4)
where TN indicates true negative and the number of pixels correctly identified as non-agricultural parcels.
F1-score is the harmonic mean of precision and recall, offering a balanced evaluation metric particularly useful under domain shift conditions or in presence of noise and uncertainty.
$F1=\frac{2\cdot Prec.\times Rec.}{Prec.+Rec.}\quad$ (5)
[…]
D.2 Object-level geometric metrics
To comprehensively evaluate the geometric quality of delineated agricultural parcels, we adopt three object-level geometric metrics: Global Over-Classification Error (GOC), Global Under-Classification Error (GUC), and Global Total Classification Error (GTC). These indicators quantify segmentation accuracy in terms of spatial overreach, omission, and overall geometric consistency.
Let $S_{i}$ denote the $i$-th predicted parcel (segmentation), and let $O_{i}$ represent the ground truth parcel that has the largest intersection area with $S_{i}$. Denote $m$ as the number of predicted parcels. The object-wise evaluation is defined as follows:
Global Over-Classification Error (GOC) measures the average extent to which predicted parcels exceed the spatial extent of their matched ground truth objects:
$\text{OC}(S_{i})=1-\frac{\text{area}(S_{i}\cap O_{i})}{\text{area}(O_{i})},$ (6)
$\text{GOC}=\sum_{i=1}^{m}\left(\text{OC}(S_{i})\cdot\frac{\text{area}(S_{i})}{\sum_{k=1}^{m}\text{area}(S_{k})}\right),$ (7)
where $\text{area}(\cdot)$ denotes the number of pixels in the respective region.
Global Under-Classification Error (GUC) quantifies the proportion of each predicted parcel not covered by the corresponding ground truth:
$\text{UC}(S_{i})=1-\frac{\text{area}(S_{i}\cap O_{i})}{\text{area}(S_{i})},$ (8)
$\text{GUC}=\sum_{i=1}^{m}\left(\text{UC}(S_{i})\cdot\frac{\text{area}(S_{i})}{\sum_{k=1}^{m}\text{area}(S_{k})}\right).$ (9)
Global Total Classification Error (GTC) synthesizes both over- and under-classification errors into one holistic metric using a root-mean-square formulation:
$\text{TC}(S_{i})=\sqrt{\frac{\text{OC}(S_{i})^{2}+\text{UC}(S_{i})^{2}}{2}},$ (10)
$\text{GTC}=\sum_{i=1}^{m}\left(\text{TC}(S_{i})\cdot\frac{\text{area}(S_{i})}{\sum_{k=1}^{m}\text{area}(S_{k})}\right).$ (11)
D.3 Edge detection evaluation metrics
For edge detection tasks on the GTPBD dataset, we evaluate model performance using three widely adopted metrics: Optimal Dataset Scale F1-score (ODS), Optimal Image Scale F1-score (OIS), and Average Precision (AP). These metrics jointly assess the accuracy, adaptability, and robustness of predicted boundaries.
Let $P_{t}$ and $R_{t}$ denote precision and recall computed at threshold $t$, and let $F_{t}$ be the corresponding F1-score:
$F_{t}=\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}.$ (12)
Optimal Dataset Scale F1-score (ODS) evaluates the global performance of an edge detector across the entire dataset using a single optimal threshold $t^{*}$:
$\text{ODS}=\max_{t\in\mathcal{T}}\left(\frac{2\cdot P_{t}^{\text{dataset}}\cdot R_{t}^{\text{dataset}}}{P_{t}^{\text{dataset}}+R_{t}^{\text{dataset}}}\right),$ (13)
where $P_{t}^{\text{dataset}}$ and $R_{t}^{\text{dataset}}$ are aggregated precision and recall over the full dataset under threshold $t$.
Optimal Image Scale F1-score (OIS) computes the mean of the per-image best F1-scores, reflecting local threshold adaptiveness:
$\text{OIS}=\frac{1}{N}\sum_{i=1}^{N}\max_{t\in\mathcal{T}}\left(\frac{2\cdot P_{t}^{(i)}\cdot R_{t}^{(i)}}{P_{t}^{(i)}+R_{t}^{(i)}}\right),$ (14)
where $P_{t}^{(i)}$ and $R_{t}^{(i)}$ denote the precision and recall on the $i$-th image under threshold $t$, and $N$ is the total number of images.
Average Precision (AP) is computed as the area under the precision–recall curve:
$\text{AP}=\int_{0}^{1}P(R),dR,$ (15)
where $P(R)$ is the precision as a function of recall, evaluated across all thresholds.
Appendix E More results
E.1 More results on Semantic Segmentation
E.1.1 Discussions the results for different regions
The radar plots in Fig. 11 compare five key metrics—Precision, Recall, IoU, Overall Accuracy, and F1-Score—across seven geographic regions for each segmentation architecture. Overall, all models achieve similar performance profiles, though U-Net demonstrates relatively average performance in terms of IoU and F1-Score, trailing behind advanced transformer-based architectures such as Mask2Former across most regions.
E.1.2 More visualization cases
Fig. 12 shows error maps for each model on the same test region, where black denotes true negatives, white true positives, blue false negatives, and red false positives. This visualization highlights that transformer-based methods such as Mask2Former [8] produce fewer false negatives in complex boundary areas. However, overall, the models struggle to accurately segment very fine-grained background details.
E.2 More visualization results on edge detection
Fig. 13 and Fig. 14 present qualitative comparisons of edge detection results on representative regions using four models: UEAD [53], PiDiNet [30], MuGE [52], and REAUNet-Sober [21]. From the visualizations, we observe that UEAD and MuGE tend to produce wider and blurred boundaries, indicating over-smoothed predictions that reduce localization precision. This issue is especially pronounced in densely terraced landscapes, where precise boundary localization is critical. In contrast, PiDiNet and REAUNet-Sober generate sharper and more compact edges, with REAUNet-Sober showing the best alignment with ground truth boundaries across complex spatial structures.
E.3 Discussions the results for different regions for terraced parcel extraction
Fig. 15 compares three parcel-extraction networks—SEANet, REAUNet, and HBGNet—over eight performance metrics (Precision, Recall, IoU, Overall Accuracy, F1-Score, GOC, GUC, and GTC) across eight geographic zones. Overall, performance variance is most pronounced in Southwest and Other International Regions, underscoring the challenge posed by highly heterogeneous terrace morphology.
E.4 Visualization cases on domain adaptation
Fig. 16, Fig. 17, and Fig. 18 illustrate the segmentation results of four UDA methods (FDA, DAFormer, HRDA, and PiPa) on six representative domain transfer tasks, with transfers targeting the North (N), South (S), and Global (G) domains, respectively. The results show that HRDA and PiPa generally produce finer boundaries and better preserve parcel structures under domain shift.
Appendix F Limitations and future work
Although GTPBD encompasses major terraced regions worldwide, the current collection is limited to six well-known terraced areas in countries where terracing is concentrated. Some atypical terraced regions in other countries are not included due to the sparse distribution of terraces in those locations. The dataset currently consists only of optical images and does not include other remote sensing modalities, such as multispectral or infrared images.
In future work, the dataset will be expanded by incorporating additional atypical terrace samples and broadening global coverage. Furthermore, GTPBD will be supplemented with multimodal geographical data, including information from Digital Elevation Models (DEMs), slope gradients, spatiotemporal features, and even will be incorporated into Q&\&&A pairs to serve multimodal large language model. Given that the dataset already covers the primary terrace types, future task may focus on domain generalization to identify new terrace types.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims are summarized in Fig. 1, with detailed explanations provided in Sec. 3 (Dataset) and Sec. 4 (Experiment).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, the limitations of this work have been discussed in Appendix F of the Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We explained our implementation details in Sec. 4.1 and hyperparameters in Appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper provides open access to the data and code, and GitHub link is offered.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The details are summarized in Sec. 4.1 and Appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Since running many repeated experiments would require significant computational resources and time, the paper reports only the direct results on our dataset.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper provides sufficient information on the computer resources in Sec. 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We affirm that our research fully complies with the NeurIPS Code of Ethics in all respects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impact in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work only investigates generalization on existing models, and poses no risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of assets used in the paper have been properly credited and the license and terms of use explicitly have been mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code and accompanying documentation in our GitHub link.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.